OpenStudy (anonymous):

What is the reciprocal of 1 1/2?

OpenStudy (anonymous):

I just can't figure this out..

jimthompson5910 (jim_thompson5910):

1 & 1/2 = 1 + 1/2 = 2/2 + 1/2 = ???

OpenStudy (anonymous):

umm I dunno I suck at math

jimthompson5910 (jim_thompson5910):

add the numerators, keep the common denominator (2) the same

OpenStudy (anonymous):

Okay..

jimthompson5910 (jim_thompson5910):

what do you get?

OpenStudy (anonymous):

None of it :\

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):